the number of radio stations on the air from 1950 to 1995 may be modeled by
f(x)214.3x-415,368, where x is the year
how man radio stations were on air in 1980
i'm not sure if i replace 1980 with x, and i'll get 8946
what is the slope of the graph
not certain
interept the slope as a rate
huh?
lol, hopefully someone can help!

Answer by adamchapman(301)

The number of radio stations at any year is
f(x) = 214.3x-415,368
in 1980, x=1980
so f(1980) = 214.3(1980)-415,368 = 8964
slope of the graph is the change in f(x) due to a unit change in x (like the slope of the line y= m*x+c is m)
so the slope is 214.3
To interpret as a rate, look at what the function f(x) represents: the number of radio stations in a year.
The rate represented by the slope is 214.3 radiostations per year.
Hope this helps,
Adam